OpenStudy (anonymous):

solve the following radical equation: equation Select one: a. Take the square root of each side to solve for x2 b. Square each side of the equation to remove the radicand c. Subtract 7 from each side d. Subtract 1 from each side

OpenStudy (anonymous):

\[\sqrt{x-1}=x-7\]

OpenStudy (anonymous):

@JoshR1213

OpenStudy (anonymous):

@SolomonZelman @satellite73

OpenStudy (anonymous):

Square both sides x-1=(x-7)(x-7) x-1=x^2-7x-7x+49 x-1=x^2-14x+49 0=x^2-14x-x+49-1 0=x^2-15x+48